solid gift wrap $4 each. printed gift wrap $6 each. the total number of gift wrap sold is 480. the total amount of money collected is $2340. how much gift wrap was sold for each kind?

Answer by rfer(10417)

x+y=480, y=480-x
4x+6y=2340
4x+6(480-x)=2340
4x+2880-6x=2340
-2x=-540
x=270 $4
y=210 $6
270*4=1080
210*6=1260